OpenStudy (anonymous):

\[\int_{-2}^{2} x~e^{-i~p~x}dx\]

OpenStudy (anonymous):

p is constant?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[u=x\]\[du=dx\]\[dv=e ^{-ipx}dx\]\[v=\frac{-e ^{-ipx}}{ip}\]\[\int\limits_{}^{}udv=uv-\int\limits_{}^{}vdu\]\[=x(\frac{-e ^{-ipx}}{ip})-\int\limits_{}^{}(\frac{-e ^{-ipx}}{ip})dx\]can you continue?

OpenStudy (anonymous):

yeah, just wanted to make sure i did it right, the first term becomes 2icos(p)/p right? and the 2nd becomes 2isin(p)/p^2 ?

OpenStudy (anonymous):

Wrote down the limits of integration on the question wrgon, they're from -1 to 1.